Let u=(-1,0,2), v=(3,1,2), w=(1,-2,-2) be vectors in standard position. Compute
1. The orthogonal complement of v
1
Expert's answer
2020-10-15T19:26:02-0400

The orthogonal compliment of "v" is

"v^{\perp}=\{ u\in V:<v,u>=0\}"

Let "u=(x,y,z)\neq 0\in v^{\perp}"

Then "<v,u>=3x+y+2z=0"

"..................(1)"

Now we have to find the nonzero solution of (1) .The free variable of equation (1) are x and z

(1) set "x=0,z=1" to obtain the solution "v_1=(0,-2,1)"

(2) set "z=0,x=1" to obtain the solution "v_2=(1,-3,0)"

The vector "v_1 \ and \ v_2" form a basis for the solution space of the equation (1) and hence a basis for "v^{\perp}."

"\therefore v^{\perp}=span\{(0,-2,1),(1,-3,0)\}"